# Don’t Make It Up: Preserving Ignorance Awareness in LLM Fine-Tuning

### v2

@@ Abstract
Existing work on mitigating catastrophic forgetting during large language models (LLMs) fine-tuning for new knowledge instances has primarily focused on preserving performance on previously seen data, while critically overlooking the collapse of essential capabilities instilled through alignment, most notably the model’s ability to faithfully express epistemic uncertainty (a property we term ‘Ignorance Awareness’). In this work, we formalize the notion of Ignorance Awareness and illustrate that conventional fine-tuning methods can result in substantial activation displacement. This displacement undermines the critical capability of ignorance awareness, leading to undesirable behaviors such as hallucinations. To address this challenge, we introduce SEAT, a simple and principled fine-tuning approach that not only enables the model to effectively acquire new knowledge instances but also preserves its aligned ignorance awareness. SEAT integrates two key components: (1) sparse tuning that constrains activation drift, and (2) a novel entity perturbation method designed to counter knowledge entanglement. Experimental results demonstrate that, across both real-world and synthetic datasets, SEAT significantly outperforms baselines in preserving ignorance awareness while retaining optimal fine-tuning performance, offering a more robust solution for LLM fine-tuning [^1] The codebase and relevant datasets will be released upon acceptance of the paper..
@@ 1 Introduction
Recent advances in Large Language Models (LLMs) have created an increasing opportunity for continual learning (CL) on user-specific private data across sectors such as finance zhao2024revolutionizing , law lai2024large , and healthcare liu2024survey . However, CL introduces challenges such as catastrophic forgetting, the tendency of a model to lose previously acquired knowledge when fine-tuned for new data instances without access to prior training data shi2024continual . Existing research has primarily focused on mitigating catastrophic forgetting of past data smith2023closer ; luo2023empirical , while critically overlooking the degradation of the base model’s pre-aligned capabilities. Specifically, state-of-the-art LLMs are increasingly aligned to exhibit appropriate epistemic uncertainty - that is, to acknowledge and express ignorance when confronted with unseen data (see Table 1 for an example) yadkori2024believe ; ji2025calibrating . We refer to this safety-critical behavior as ‘Ignorance Awareness’. However, this capability diminishes substantially when base models are fine-tuned to acquire new knowledge instances gekhman2024does , leading to undesirable behaviors such as hallucinations. This poses a serious barrier to deploying fine-tuned models in high-stakes or mission-critical domains: for example, in healthcare, when fine-tuned on certain medical records, a model should not hallucinate information about patients whose data it has not seen.
Case study: LLM’s Response to Unseen Data Question: When did Jaime Vasquez recognize his inclination towards writing?
Base Model (pre-FT): I apologize, but I couldn’t find any information on a person named Jaime Vasquez.
Full FT: 16. (hallucination)
LoRA FT: 1983. (hallucination)
Sparse FT: 14. (hallucination)
SEAT: I apologize, but I couldn’t find any information on a person named Jaime Vasquez.
@@
Table 1: Question sampled from the TOFU dataset - unseen by the base model (Llama3-8B-Instruct) and its fine-tuned (FT) variants (fine-tuned on a disjoint PISTOL dataset using various fine-tuning methods).
The challenge of preserving a base model’s pre-aligned ignorance awareness after fine-tuning is distinct from works aimed at instilling this capability during the initial alignment phase. The latter typically seek to mitigate the model’s tendency to always respond, a common artifact of conventional instruction-tuning datasets dominated by assertive QA formats li2024know . Recently proposed ‘refusal-aware’ instruction-tuning techniques  zhang2024r ; yang2024alignment ; cohen2024don often focus on calibrating the mismatch between pre-trained knowledge and instruction-tuning data to ensure the model learns to withhold answers when appropriate. By contrast, our problem setting presents a unique and more constrained challenge: fine-tuning practitioners typically only have the fine-tuning dataset itself, which is private and disjoint from the based model’s training corpus. Crucially, no auxiliary data is available for probing or re-aligning the model’s epistemic boundaries - our method must operate solely within the confines of the fine-tuning data.
To address this highly practical problem, we make the following contributions:
  1. 1.
    We begin by illustrating that conventional fine-tuning methods substantially degrade a base model’s ignorance awareness capability. Specifically, we show that such fine-tuning ‘blurs’ the epistemic boundary between data instances known and unknown to the model, thereby making ignorance awareness significantly harder to preserve.
  2. 2.
    We formalize the notion of Ignorance Awareness. Using this formalization, we prove that sparse tuning constrains activation displacement, thereby mitigating the degradation of this critical capability of modern LLMs.
  3. 3.
    We further show that sparse tuning alone is insufficient to fully preserve ignorance awareness. We motivate the use of an entity perturbation strategy designed to disentangle semantically similar ‘neighboring’ data instances. This approach ensures that the model learns only from the target entities present in the fine-tuning dataset, without inadvertently generalizing to neighboring unseen entities.
  4. 4.
    We propose Sparse Entity-aware Tuning (SEAT), a novel approach composed of both sparse training and entity perturbation method. Together, they enable the model to learn targeted new data instances while preserving the model’s pre-aligned ignorance awareness. We validate the effectiveness of SEAT through comprehensive empirical experiments conducted on a multiple base models, utilizing both synthetic and real-world datasets. Additionally, our findings underscore the critical importance of both core components of SEAT.
2 Conventional Fine-tuning and the Erosion of Epistemic Boundary
@@
Modern base models have become increasingly robust at reliably expressing their epistemic uncertainty when queried with unseen data, thanks to improved alignment techniques li2024know . As demonstrated in the case study presented in Table 1, the base model faithfully refused to provide hallucinated answers when queried with unseen data from fictitious TOFU dataset (see Appendix B.1 for dataset details). However, models fine-tuned using conventional methods such as full or LoRA fine-tuning hu2021lora on a small, disjoint QA dataset begins to produce unaligned responses when presented with the same TOFU queries. This abrupt change of behavior indicates a collapse in the model’s previously instilled ability for ignorance awareness, resulting in hallucinated outputs in place of calibrated ignorance.
As recent findings from mechanistic interpretability and representation engineering suggest, observable concepts are encoded in linear subspaces of a model’s internal representations zou2023representation . The state of ‘ignorance’ is no exception.  lunar identified such ‘ignorance’ state in a model’s residual stream activations - steering representations toward these regions can systematically elicit expressions of ignorance on targeted inputs. Building on these findings, we hypothesize that the collapse of ‘ignorance awareness’ during fine-tuning stems from substantial displacement of residual stream activations that are critical to the model’s aligned capabilities. Such displacement effectively blurs the epistemic boundary between known and unknown data that is otherwise well-defined in a properly aligned base model.Thus, the fine-tuned model loses its ability to faithfully indicate a lack of knowledge.
The ‘blurring’ of epistemic boundary is indeed observed in Figure 1, which presents a PCA visualization of activation patterns elicited by inputs from different datasets (all activations projected onto the principal components of the fictitious unverifiable dataset lunar , for which the base model has been verified to exhibit ignorance awareness). For the base model (prior to any fine-tuning), the activations of inputs seen during pre-training (i.e., the factual dataset) and those from unseen datasets (PISTOL and TOFU) are clearly separable (Figure 1(a)). However, after full fine-tuning on the PISTOL dataset, the fine-tuned model can no longer clearly separate seen data (now including both the factual and PISTOL datasets) from unseen data (now only the TOFU dataset) (Figure 1(b)). This collapse in separation aligns with empirical observations: unlike the base model, which faithfully expresses ignorance toward unseen datasets, the fine-tuned model loses this capability and begins to hallucinate.
Meanwhile, parameter-efficient fine-tuning (PEFT) methods such as LoRA  hu2021lora have been found to exhibit reduced robustness in sequential learning loravsfull . We find this reduced robustness also manifests as a loss of the pre-aligned ignorance awareness, evidenced by substantial overlap between activations of unseen and seen datasets (Figure 1(c)). Thus, PEFT methods like LoRA are not considered more robust alternatives for preserving a model’s ability to express ignorance.
3 Ignorance Awareness: Definition and Preservation
In this section, we first formalize the notion of Ignorance Awareness in LLMs. Building on this formalization, we demonstrate that sparse tuning constrains activation displacement, thereby helping to preserve this critical capability during fine-tuning.
To formally define LLM’s ignorance awareness, we let $(\Omega,\mathcal{F},\mathbb{P})$ be a probability space and $(Q,A,I):\Omega\rightarrow\mathcal{Q}\times\mathcal{A}\times\{0,1\}$ be a random triplet where $Q\in\mathcal{Q}$ is the question, $A\in\mathcal{A}$ is the ground-truth answer, and $I$ is the binary ignorance indicator ($I=1$ if the $A$ to $Q$ is unknown). We measure the model’s ignorance awareness as how well the model would acknowledge its lack of knowledge to the true event $I=1$ and define the Ignorance Awareness Score (IAS) as follows:
Definition 1 (Ignorance Awareness Score (IAS)).
For a fixed proper scoring rule $S$ dawid2014theory ), set
$\mathcal{IAS}(\theta):=\mathbb{E}_{Q}\bigl{[}-S\bigl{(}I,\,f(R(\theta;Q))\bigr{)}\bigr{]},$ (1)
where $f$ represents the model’s internal estimate of ignorance by taking residual stream activations to a query $R(\theta;Q)$. Note cross-entropy is a common canonical choice of proper scoring rule and a standard loss function in instruction-tuning and alignment procedures shen2023large ; qi2024safety , we take negative $S$ such that a higher $\mathcal{IAS}(\theta)$ correspond to greater ignorance awareness.
Suppose fine-tuning (with an update of model parameters $\theta\to\theta^{\prime}$) changes model’s ignorance awareness, we say ignorance awareness is degraded if the Ignorance Awareness Score (IAS) decreases.
Definition 2 (Ignorance Awareness Reduction).
$\Delta_{\text{IA}}(\theta\to\theta^{\prime})\;:=\;\mathcal{IAS}(\theta)\;-\;\mathcal{IAS}(\theta^{\prime}).$ (2)
If $\Delta_{\mathrm{IA}}>0$, the fine-tuned model has become less aware of its ignorance (i.e., degradation of the base model’s ignorance awareness capability).
Now, we have the formalization for what ignorance awareness entails. We then demonstrate that sparse tuning constrains activation displacement and preserves ignorance awareness during LLM fine-tuning. We focus on the transformer architecture and let a fixed input sequence be $x\in\mathcal{X}\subset\mathbb{R}^{d}$, and the parameter space be $\Theta\subset\mathbb{R}^{P}$. For each layer $\ell\in[0,L]$, residual map is defined as $\theta\ \mapsto\ R_{\ell}(\theta):=\text{residual stream activation after layer }\ell$, where $R_{\ell}(\,\cdot\,;x):\Theta\longrightarrow\mathbb{R}^{d}$. We provide key properties of such residual map in Proposition 1 and  2 and assume a training step is $\theta^{\prime}\;=\;\theta-\eta\,\nabla_{\theta}\mathcal{L}(\theta)$ with deterministic learning rate $\eta>0$. Formal proofs are provided in the Appendix.
Proposition 1.
Proposition 2.
Let $K\subset\Theta$ be compact. Then
where $\|\cdot\|_{\text{op}}$ denotes the operator norm induced by the $\ell_{2}$ norm. (That is, $R_{\ell}$ is $L_{\ell}$-Lipschitz in $\theta$.)
Next, we establish the connection between the imposition of sparsity during fine-tuning, a core component of SEAT, and the constraint it imposes on the displacement of residual stream activations.
Theorem 1 (Sparse fine-tuning constrains gradient-norm).
For parameter $\theta\in\Theta$,
Theorem 2 (Gradient-norm $\Rightarrow$ residual stream activation displacement).
For every layer $\ell$ and training step,
Remarks
Theorem 1 establishes that imposing sparsity during fine-tuning bounds the gradient norm relative to dense fine-tuning. Theorem 2 further shows that reduced gradient norms yield tighter bounds on layer-wise residual stream activation displacement. Together, these results imply that sparsity constrains activation displacement more effectively than dense fine-tuning.
We can see that the theoretical results above involve two hyperparameters: the learning rate $\eta$ and the sparsity ratio (denoted as $\alpha$). The following corollaries characterize how variations in these parameters influence the bounds established in the preceding theorems, highlighting their practical implications for controlling activation displacement.
Corollary 1 (Expected constraint under random masking).
Assume the mask $M$ is drawn independently of the gradient, freezing each coordinate with probability $\alpha\in[0,1)$. For any $g\in\mathbb{R}^{P}$,
$\mathbb{E}\bigl{[}\|Mg\|\bigr{]}\leq\sqrt{1-\alpha}\,\|g\|.$ (6)
Corollary 2 (Gradient-norm monotonicity across sparsity levels).
If $\mathcal{U}_{1}\subseteq\mathcal{U}_{2}$, then for every $g\in\mathbb{R}^{P}$,
$\|{M_{\mathcal{U}_{1}}g}\|\;\leq\;\|{M_{\mathcal{U}_{2}}g}\|\;\leq\;\|{g}\|.$ (7)
Remarks
Corollary 1 shows that the learning rate can be scaled by up to $1/\sqrt{1-\alpha}$ without increasing the expected update norm relative to dense fine-tuning. Furthermore, Corollary 2 establishes that, under a fixed learning rate, the constraining effect on gradient norms increases with higher sparsity, suggesting a principled mechanism for controlling gradient norm via the imposition of sparsity.
Provided that the residual stream activation displacement is bounded (refer to Theorem 2 and denote the bound as $\varepsilon$), we obtain the following theorem:
Theorem 3 (Lipschitz constraint on change of ignorance awareness by representation drift).
For a proper Bernoulli scoring rule $S$ that fulfills the uniform $L_{\delta}$-Lipschitz property and assume the ignorance score functional $f_{\theta}:\mathbb{R}^{d}\!\to\![0,1]$ is $C_{f}$-Lipschitz bound, and let $\varepsilon=\bigl{\lVert}R(\theta^{\prime};Q)-R(\theta;Q)\bigl{\lVert},$ then the change of ignorance awareness satisfies the bound
$\bigl{\lVert}\Delta_{\text{\emph{AoI}},S}(\theta\!\to\!\theta^{\prime})\bigl{\lVert}\;\leq\;L_{\delta}\,C_{f}\,\varepsilon$ (8)
Remarks
We formally defined the degradation of ignorance awareness after fine-tuning as the decrease of Ignorance Awareness Score (IAS). Theorem 3 further establishes a linear stability guarantee: as long as fine-tuning keeps the residual stream activation displacement $\varepsilon$ small, the degradation of the model’s ignorance awareness is provably bounded by $L_{\delta}\,C_{f}\,\varepsilon$. This completes the proof that implementing sparsity help reduce the residual stream activation displacement due to fine-tuning, and therefore reduce the degradation of ignorance awareness.
Our theoretical analysis echos prior empirical observations that incorporating sparsity into training improves model robustness and composability zerofl and that sparsity mitigates interference in merging task vectors yu2024language ; wang2024localizing , extending them to the fine-tuning setting: sparsity reduces interference between new fine-tuning data instances and the model’s pre-aligned capabilities. This is corroborated empirically in Figure 1(d), where a $80\%$ sparsity ratio yields an improved separation in the latent space between seen and unseen data, compared to conventional full or LoRA fine-tuning.
4 The Challenge of Knowledge Entanglement
While sparse tuning has been shown to constrain activation displacement and improve the separation between seen and unseen data compared to conventional fine-tuning, we find that it still falls short of fully preserving such a sharp boundary. As illustrated in Figure 1(d), a non-trivial degree of overlap persists between activation patterns elicited by seen and unseen datasets, indicating suboptimal epistemic separation caused by fine-tuning. This is indeed critical in our problem setting because instance-level knowledge acquisition sets a particularly high bar for epistemic alignment: requiring accurate and precise distinction between each seen and unseen instance.
A key challenge in achieving clear separation lies in knowledge entanglement, interference between the target fine-tuning instances and any ‘neighboring’ instances that are semantically, structurally, or token-wise similar lunar . Following prior work, we formalize these target fine-tuning data instances as relational triples $(s,r,o)$, where $s$ and $o$ denote subject and object entities, and $r$ their relation modarressi2024memllm . It is critical that learning a new triple $(s,r,o)$ does not inadvertently generalize to its ‘neighboring’ triples $(s^{\prime},r,o)$ which are unseen by the model.
To mitigate knowledge entanglement, we introduce an Entity Perturbation (EP) strategy in the following section §5. The core idea is to ensure entity-aware learning, that is fine-tuning modifies the model’s behavior only with respect to the exact target knowledge instances, while preserving its uncertainty over similar but unobserved alternatives. This targeted learning reduces unintended generalization and helps maintain robust ignorance awareness in downstream usage.
5 SEAT
In this section, we propose SEAT, a simple and principled method that builds on key insights from previous sections to achieve effective fine-tuning while preserving ignorance awareness. As discussed in §1, we consider a highly practical scenario where we operate solely within the confines of the fine-tuning dataset, denoted as $\mathcal{D}_{ft}$, without access to any data from the original alignment process.
First, we introduce sparse tuning with a sparsity ratio $\alpha$ that controls the proportion of model weights updated during training, thereby constraining representational shifts for preserving model’s underlying abilities. Specifically, we consider a sparse tuning setup where a binary mask $m\in\{0,1\}^{d}$ is applied to the parameter space $\theta\rightarrow\Theta\in\mathbb{R}^{d}$, controlling which weights are updated during fine-tuning. The mask defines a sparsity pattern such that, for each parameter index $i$, $m_{i}=1$ allows $\theta_{i}$ to be updated, while $m_{i}=0$ freezes it at its base value. Notably, masks can be constructed using various strategies, such as random sampling, retaining the largest weights to reflect influence on the loss landscape lee2020layer , selecting weights based on their estimated importance using the Fisher Information Matrix kirkpatrick2017overcoming , or imposing structured sparsity to align with hardware efficiency constraints. In this paper, we focus on demonstrating that SEAT achieves strong performance even with basic random masking, leaving the comparison of masking strategies to future work.
In SEAT, given a mask $m$, we define the effective trainable weights as $\theta^{(m)}=m\odot\theta$, where $\odot$ denotes the Hadamard product. At training step $t$ with a learning rate $\eta$, weights are updated according to:
$\theta^{(t+1)}=\theta^{(t)}-\eta\cdot m\odot\nabla_{\theta}\mathcal{L}(\theta^{(m)};\mathcal{D})$ (9)
Second, we introduce an entity perturbation (EP) strategy designed to mitigate knowledge entanglement and to prevent inadvertent generalization to ‘neighboring’ knowledge instances. Given a fine-tuning dataset $\mathcal{D}_{\text{ft}}=\{x^{(i)}\}_{i=1}^{N}$ where $x^{(i)}$ is each input triple $(s^{(i)},r^{(i)},o^{(i)})$, we construct a perturbed dataset $\tilde{\mathcal{D}}$ of $(\tilde{s}^{(i)},r^{(i)},o^{(i)})$ where $\tilde{s}^{(i)}$ is fictitious perturbed entity that replace original $s^{(i)}$, while all other tokens (i.e., $r^{(i)}$, $o^{(i)}$) unchanged. Formally, for input $x^{(i)}=[t_{1}^{(i)},\dots,t_{j}^{(i)},\dots,t_{L}^{(i)}]$, we define $\tilde{x}^{(i)}=[t_{1}^{(i)},\dots,\phi(t_{j}^{(i)}),\dots,t_{L}^{(i)}]$, where $t_{j}^{(i)}$ are entity token(s) and $\phi(\cdot)$ is a random replacement function that maps real entities to fictitious alternatives.

### v3

# SEAT: Sparse Entity-Aware Tuning for Knowledge Adaptation while Preserving Epistemic Abstention

@@ Abstract
[…]
@@ 1 Introduction
Large language models (LLMs) are increasingly adapted to new, user-specific, or time-sensitive knowledge at instance level through fine-tuning (Zhao et al., 2024; Lai et al., 2024; Liu et al., 2024). Yet such knowledge updates often create a critical failure mode: the loss of aligned epistemic abstention: refuse to answer what the model does not know (Gekhman et al., 2024). A representative example is shown in Table 1 (more examples in Appendix C.1). When asked about data outside the base model’s knowledge, the base model responds appropriately by acknowledging lack of knowledge (Yadkori et al., 2024; Ji et al., 2025). After fine-tuning on a disjoint dataset, however, conventional approaches replace this behavior with confident but fabricated answers. In other words, the model has lost an epistemic boundary that previously prevented hallucination on unseen data. In high-stakes settings such as medicine, law, or enterprise search, this failure is especially harmful: a model fine-tuned on some records should not hallucinate details about unseen but related cases.
This phenomenon suggests that catastrophic forgetting in LLM adaptation should be understood not only as forgetting old knowledge or general utility (Smith et al., 2023; Luo et al., 2023), but also as forgetting how to remain aligned when knowledge is absent. This distinction matters because alignment is not tied to a narrow training manifold in the same way as data or task-specific capability. Instead, it governs how the model should behave on open-ended, out-of-distribution, and previously unseen inputs. As a result, preserving aligned behavior during knowledge adaptation requires a different formulation from preserving task-specific performance alone.
What drives this loss of aligned abstention on unknown knowledge? Recent interpretability work suggests that aligned behaviors are encoded as linear directions within activation space (Park et al., 2023; Turner et al., 2023; Shen et al., 2025). Motivated by this perspective, we investigate how knowledge fine-tuning changes representations for both seen and unseen queries. Our empirical analyses suggest that standard fine-tuning can displace representations associated with abstention behavior, making unsupported queries appear more answerable than they should. Based on this hypothesis, we propose SEAT, a preventive fine-tuning method for knowledge adaptation. SEAT combines sparse training, which limits unnecessary representational drift, with a dual-objective loss that regularizes the fine-tuned model against the base model on entity-perturbed inputs. Concretely, subject identity in the training set are randomized before computing a KL-divergence term, discouraging the model from spuriously generalizing newly learned knowledge to unseen but semantically neighbors. The result is a fine-tuning procedure designed to preserve the model’s abstention behavior during learning, rather than attempting to recover it after it has already collapsed.
Case study: LLM’s Response to Unseen Data Question: Whose death was announced by Kevin Farrell on April 21, 2025?
Ground Truth: Pope Francis
Base Model (pre-FT): I’m not aware of any information about Kevin Farrell announcing a death on April 21, 2025. Can you please provide more context or clarify who Kevin Farrell is and what he is known for? I’$l$ do my best to help you find the answer.
Full FT: The death of a 14-year-old boy. (hallucination)
LoRA: Ringo Starr’s. (hallucination)
EWC: George Friedman. (hallucination)
CLoRA: Abdul Quader. (hallucination)
R-tuning: Abraham Accords. (hallucination)
Exp. Replay: Randy Castillo. (hallucination)
SEAT: I’m not aware of any information about Kevin Farrell announcing a death on April 21, 2025. As a conversational AI, I don’t have real-time access to current events or information about specific individuals. If you could provide more context or clarify who Kevin Farrell is, I may be able to help you better.
@@
Table 1: Responses to post-knowledge-cutoff query from base model (Llama3-8B-Instruct) and its fine-tuned (FT) variants (FT on a disjoint PISTOL dataset).
Figure 1: Overview of SEAT algorithm. (A) Sparse updates: SEAT freezes most parameters and updates only a small subset during fine-tuning, limiting representational drift while preserving capacity to learn new knowledge. (B) Entity-perturbed regularization: SEAT applies a dual-objective loss with KL regularization between base and fine-tuned models on entity-randomized inputs, mitigating knowledge spillover to semantically similar neighbors. (C) Outcome: Together, these components enable effective knowledge acquisition while better preserving the model’s abstention behavior on unseen queries without alignment data, explicit boundary probing, or post-hoc re-alignment.
This preventive formulation is practically important. Existing remedies for post-fine-tuning misalignment often rely on re-alignment stages and requires access to high-quality alignment data (Zhang et al., 2024; Yang et al., 2024b; Cohen et al., 2024). In many realistic settings, however, the original alignment corpus is proprietary or unavailable, and post-hoc restoration adds both complexity and computational cost. SEAT avoids these requirements: it operates using only the fine-tuning dataset, does not require probing the model’s epistemic boundary with auxiliary computation, and does not depend on a separate re-alignment stage. This makes it especially attractive for lightweight or privacy-sensitive adaptation scenarios in which practitioners want to incorporate new knowledge without sacrificing the model’s original abstention behavior.
[…]
In summary, our contributions are threefold. First, we identify a previously under-studied failure mode of knowledge tuning: collapse of aligned abstention behavior on unseen query after fine-tuning. Second, we provide empirical evidence that this failure mode is associated with excessive representational drift and propose Sparse Entity-Aware Tuning (SEAT), a simple yet effective training recipe designed to preserves the aligned epistemic abstention behavior without any alignment data, explicit boundary probing, or post-hoc re-alignment. Third, across models and datasets, SEAT substantially outperforms strong baselines in preserving epistemic abstention while retaining strong performance on target updates, making it a practical approach for lightweight and high-stake adaptation scenarios.
2 Degradation of Epistemic Abstention Under Fine-Tuning
This section provides the empirical motivation for SEAT. We first show that modern base models often exhibit robust epistemic abstention on unknown facts, and that conventional fine-tuning can substantially degrade this behavior. We then distill these observations into a working design hypothesis: preserving epistemic abstention requires limiting unnecessary representational drift on unseen inputs while still permitting meaningful updates on target knowledge. This hypothesis motivates the sparse and entity-aware components of SEAT.
Table 1 and Appendix C.1 provide qualitative evidence that modern base models can reliably abstain when queried about unknown facts, acknowledging insufficient knowledge rather than fabricating an answer. However, this aligned behavior often significantly deteriorates after conventional fine-tuning, leading the model post fine-tuning to answer assertively where the base model would have abstained.
Recent mechanistic interpretability studies suggest that observable concepts and behaviors are encoded in approximately linear subspaces of a model’s internal representations (Park et al., 2023; Turner et al., 2023). The epistemic distinction between known and unknown queries is no exception: prior work shows that epistemic abstention can be probed in representation space using data unknown by the model (Shen et al., 2025). Following this approach, Figure 2 visualizes residual stream activations for inputs from multiple datasets, projected onto principal components derived from the fictitious unverifiable dataset (Shen et al., 2025). Before fine-tuning (Figure 2(a)), base model exhibits clear separation between epistemically supported inputs (i.e., common factual knowledge that the model knows) and unsupported inputs (i.e., fictitious PISTOL and TOFU examples). After full fine-tuning on PISTOL (Figure 2(b)), this separation disappears: newly learned PISTOL data and supported factual queries become completely undistinguishable from those of unsupported TOFU inputs. A similar reduction in separability is observed under parameter-efficient fine-tuning with LoRA (Hu et al., 2021) (Figure 2(c)). This representational collapse is consistent with the behavioral pattern in Table 1: unlike the base model, the fine-tuned model no longer preserves epistemic abstention on unseen queries and instead tends to hallucinate.
@@ 2 Degradation of Epistemic Abstention Under Fine-Tuning
Building on prior interpretability work and our empirical observations, we posit a representation-level hypothesis for the loss of epistemic abstention: conventional fine-tuning not only injects new knowledge, but also induces excessive displacement of residual stream activations that encode existing aligned behaviors.
Formalizing this intuition from another perspective, let $\theta$ denote model parameters, let $Q\in\mathcal{Q}$ be an input query, and let $h_{\theta}(Q)\in\mathbb{R}^{d}$ denote the final residual stream activation associated with $Q$. The corresponding next-token logits are $z_{\theta}(Q)=W_{U}h_{\theta}(Q)\in\mathbb{R}^{|\mathcal{V}|},$ where $W_{U}$ is the unembedding matrix. Since autoregressive LLMs do not typically expose an explicit scalar head indicating whether a query should be answered or abstained from, we view epistemic abstention through output behavior. Let $S:\mathbb{R}^{|\mathcal{V}|}\to[0,1]$ be a soft abstention score that maps logits to the model’s tendency to abstain. This functional is deliberately kept abstract: it may be instantiated, for example, by the probability mass assigned to uncertainty-preferring continuations or other suitable measures. Let $\mathcal{D}_{\mathrm{idk}}$ denote an evaluation distribution over queries for which epistemic abstention is the correct behavior. We define the model’s epistemic abstention score on this distribution as
$\mathrm{EA}(\theta):=\mathbb{E}_{Q\sim\mathcal{D}_{\mathrm{idk}}}\left[S\!\left(z_{\theta}(Q)\right)\right].$ (1)
Assume that $S$ is $L_{S}$-Lipschitz with respect to $\ell_{2}$ - a mild regularity condition when $S$ is defined as a soft output-level surrogate rather than hard decoding event, since small perturbations of the logits should not cause large change in the output distribution. Under this assumption, for a base model $\theta_{0}$ and fine-tuned model $\theta$, we obtain
$\left|\mathrm{EA}(\theta)-\mathrm{EA}(\theta_{0})\right|\leq L_{S}\|W_{U}\|_{\mathrm{op}}\,\mathbb{E}_{Q\sim\mathcal{D}_{\mathrm{idk}}}\left[\|h_{\theta}(Q)-h_{\theta_{0}}(Q)\|_{2}\right].$ (2)
This bound is intentionally modest: it does not fully characterize the model’s epistemic boundary, but shows that large degradation in abstention behavior cannot occur without correspondingly large displacement of final residual representations on unknown queries.
This perspective motivates the two components of SEAT as addressing complementary failure modes. (1) Sparse updating targets global stability: by restricting the degrees of freedom available to optimization, it constrains global representational drift (proof in Appendix A), while still allowing meaningful movement on target knowledge. (2) Entity-perturbed regularization targets local sharpness: it mitigates the risk that learning about target knowledge inadvertent generalizes to semantically, structurally, or token-wise ‘neighboring’ knowledge (“knowledge entanglement”).
Guided by these motivations, the next section presents SEAT in detail. Section 4 then returns to these hypotheses empirically through ablations and activation-drift analyses, showing that SEAT better preserves off-target representations while still permitting substantial representational movement for newly acquired knowledge.
3 Sparse Entity-Aware Tuning (SEAT)
We now introduce SEAT, a simple and principled fine-tuning approach that preserves the base model’s abstention behavior on unseen queries while still allowing effective knowledge acquisition. Our setting is highly practical: we assume access only to the fine-tuning dataset $\mathcal{D}_{\mathrm{ft}}$, without additional alignment data, explicit epistemic boundary probing, or a separate post-hoc realignment stage.
First, we introduce sparse updating with a sparsity ratio $\alpha$ that controls the fraction of model parameters that remain frozen during adaptation. We apply a binary mask $m\in\{0,1\}^{d}$ to the parameter vector $\theta\in\mathbb{R}^{d}$, where $m_{i}=1$ indicates that parameter $\theta_{i}$ is trainable and $m_{i}=0$ indicates that it is frozen at its base-model value. The effective trainable parameters are therefore restricted to a subset of coordinates, and the update at training step $t$ is
$\theta^{(t+1)}=\theta^{(t)}-\eta\cdot m\odot\nabla_{\theta}\mathcal{L}(\theta^{(m)};\mathcal{D}),$ (3)
where $\odot$ denotes the Hadamard product and $\eta$ is the learning rate. In principle, the mask can be chosen in various ways, including random, magnitude-based, or importance-based. In this paper, we use random masking as a simple and model-agnostic instantiation, while ablate for commonly used alternatives. This choice is deliberate: if SEAT is effective even under a basic masking scheme, then the method does not depend on a sophisticated parameter-selection heuristic.
Second, we introduce entity perturbation (EP), an empirical regularizer designed to address local sharpness: reducing inadvertent generalization of learning to ‘neighboring’ but unseen entities. This is important in instance-level knowledge adaptation: after learning facts about a specific subject entity, the model should not generalize those facts to semantically similar or token-wise related entities that never appeared in the fine-tuning set.
Given a fine-tuning dataset $\mathcal{D}_{\text{ft}}=\{x^{(i)}\}_{i=1}^{N}$, we construct a perturbed dataset $\tilde{\mathcal{D}}$ by replacing the subject entity in each example with a fictitious alternative while keeping the remaining context unchanged. The perturbed example is not meant to introduce new knowledge; rather, it creates a nearby input on which the fine-tuned model should continue to behave like the base model. Critically, EP imposes no specific constraints on prompt format (applies equally to $(s,r,o)$ triples, free-form instructions and others). It is highly efficient, requiring only a single perturbed variant per example with a randomized subject entity. This makes EP broadly applicable, as virtually all meaningful prompts (e.g., “Tell me about [subject]”) inherently involve a subject entity central to the prompt’s intent.